Find
a)-A raised to -1+3B raised to T
b)B raised to -1+(A raised to T+A raised to -1)
1
Expert's answer
2020-05-11T12:29:43-0400


Find:

a)-A raised to -1+3B raised to T


"let\\A=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}\\B=\begin{bmatrix}
 i & j\\
 k & l
\end{bmatrix}"

then

"-A=\begin{bmatrix}
 -a & -b \\
 -c & -d
\end{bmatrix} \implies -A^{-1}=\frac{1}{ad-cb}\begin{bmatrix}
 -d& b \\
 c &- a
\end{bmatrix}"


"B^{T}=\begin{bmatrix}
 i & k\\
 j& l
\end{bmatrix} \implies 3B^{T}=\begin{bmatrix}
 3 i & 3k\\
 3 j& 3l
\end{bmatrix}"


"-A^{-1}+3B^T=\frac{1}{ad-cb}\begin{bmatrix}
 -d& b \\
 c &- a
\end{bmatrix}+\begin{bmatrix}
 3 i & 3k\\
 3 j& 3l
\end{bmatrix}"


"-A^{-1}+3B^T=\begin{bmatrix}
 \frac{3i(ad-cd)-d}{ad-cd} & \frac{3k(ad-cd)+b}{ad-cd}\\
 \frac{3j(ad-cd)+c}{ad-cd} & \frac{3l(ad-cd)-a}{ad-cd}
\end{bmatrix}"


b)B raised to -1+(A raised to T+A raised to -1)


"B^{-1}=\frac{1}{il-kj}\begin{bmatrix}
 l&- j\\
 - k &i
\end{bmatrix};\phantom{i}A^{T}=\begin{bmatrix}
 a& c\\
 b &d
\end{bmatrix};A^{-1}=\frac{1}{ad-cb}\begin{bmatrix}
 d& -b \\
 -c &a
\end{bmatrix}"


"B^{-1}+A^T+A^{-1}=\begin{bmatrix}
 \frac{l}{il-kj}+a+\frac{d}{ad-cb}& \frac{j}{kj-il}+c+\frac{b}{cb-ad}\\
 \frac{k}{kj-il}+b+\frac{c}{cb-ad} & \frac{i}{il-kj}+d+\frac{a}{ad-cb}
\end{bmatrix}"